Solve the following equations using Wronskian’s method d³y/dx³+3d²y/dx²+3dy/dx+y=(x+40)e^-x


1
Expert's answer
2021-02-24T07:07:45-0500

y"'+3y"+3y'+y=(x+40)e-x

Let x=t

Let the general solution be

"y(t) =c_1y_1(t) +c_2y_2(t)+ c_3y_3(t)+y_p(t)"


y1, y2 , y3-are the solutions of the homogeneous equation. These solutions may be obtained by the characteristic equation

"\lambda^3+3\lambda^2+3\lambda+1=0"

"(\lambda+1)^3=0"

"\lambda_1=\lambda_2=\lambda_3=-1"

"y_1=y_2=y_3=e^{-t}"

Wronskian test for independence.

Let us calculate the Wronskian:

"W=det\begin{pmatrix}
 y_1 & y_2 & y_3 \\
 y_1' & y_2' & y_3' \\
y_1'' & y_2'' & y_3''
\end{pmatrix}=det\begin{pmatrix}
 e^{-t} & e^{-t} & e^{-t} \\
 -e^{-t} & -e^{-t} & -e^{-t} \\
e^{-t} & e^{-t} & e^{-t} 
\end{pmatrix}=3e^{-t}(-e^{-2t}+e^{-2t})=0"

W=0, so y1, y2,y3 are dependent. So

"y_p(t)=f(t)=(x+40)e^{-t}"

Answer: y(x)=3ce-x+(x+40)e-x=e-x(3c+x+40)